OpenStudy (anonymous):

Prove 16sin10sin30sin50sin70sin90=1

OpenStudy (anonymous):

sin 90=1 and sin 30=1/2 putting them you will get 8 sin 10 sin 50 sin 70 =1 :)

OpenStudy (anonymous):

And how am I supposed to continue?

OpenStudy (anonymous):

now 2sin 10 sin 50=???????

OpenStudy (anonymous):

@ErkoT

OpenStudy (anonymous):

That, my dear Sir, is an excellent question.

OpenStudy (anonymous):

2 sin a sin b=cos(a-b)-cos(a+b) put it there

OpenStudy (anonymous):

4[cos40-cos 60]sin 70 now cos 60=1/2 2[2cos40-1]sin70 what is the value of 2cos2x-1=????????

OpenStudy (anonymous):

@ErkoT

OpenStudy (anonymous):

@Yahoo! do it..........!!!!!!!!!

OpenStudy (anonymous):

cos^2x-3sin^2x?